OpenStudy (anonymous):

a sled is initially given a shove up a frictionless 28.0 degree incline.it reaches a maxinum vertical height 1.35m higher than where it started. what was the initial speed? if the incline has friction of f=20N and the sled has mass m=5kg, determine the vertical height reached by the sled suppose it is given the same initial speed as in 5.14m/s.

OpenStudy (aaronq):

Use the conservation of energy for the first part: \(\sf KE_f+PE_f=KE_i+PE_i\) For the second part you need to sum up the forces and use newton's second law  \(\sf \sum F_x=-sin(28^o)F_g-f=ma\) \(\sf \sum F_x=n-sin(28^o)F_g=0\) Use \(\sf v_f^2=v_i^2+2a\Delta x\) find the distance travelled up the ramp Use trigonometry to find the height reached.